OpenStudy (chosenmatt):

What is the approximate area of a circle with a diameter of 46 meters?

OpenStudy (chosenmatt):

@iGreen

OpenStudy (igreen):

A = pi * r^2 The radius is half the diameter. 46 / 2 = ?

OpenStudy (chosenmatt):

23

OpenStudy (igreen):

Yes, so the radius is 23. A = 3.14 * 23^2

OpenStudy (igreen):

Multiply 3.14 * 23 * 23

OpenStudy (chosenmatt):

1661.06

OpenStudy (igreen):

Yes, that's our answer.